# Differentially Private Optimization with Sparse Gradients

### v1

@@ 4 Lower Bounds for DP Mean Estimation with Sparse Data
We provide matching lower bounds to the upper bounds obtained in the previous section. Note that although the lower bounds we state are for mean estimation, known reductions imply analogous lower bounds for DP-ERM and DP-SCO [6, 7]. First, for pure-DP we provide a new packing-type construction based on sparse vectors. This construction is used later in a block-diagonal form, which provides the right low/high-dimensional transition. On the other hand, for approximate-DP, a block diagonal reduction with existing fingerprinting codes [12, 34], suffices to obtain lower bounds that exhibit a nearly tight low/high-dimensional transition. For simplicity, we consider the case of $L=1$, i.e. ${\cal S}_{s}^{d}=\{z\in\mathbb{R}^{d}:\,\|z\|_{0}\leq s,\|z\|_{2}\leq 1\}$; it is easy to see that any lower bound scales linearly in $L$.
@@ 4.1 Lower Bounds for Pure-DP
Given a sample space ${\cal Z}$, we say that a probability $\mu$ over ${\cal Z}^{n}$ induces an $(\alpha,\beta)$-distributional lower bound for $(\varepsilon,\delta)$-DP mean estimation if
$\inf_{{\cal A}:\,(\varepsilon,\delta)\text{-DP}}\mathbb{P}_{S\sim\mu,{\cal A}}\Big{[}\|{\cal A}(S)-\bar{z}(S)\|_{2}\geq\alpha\Big{]}\geq\beta.$
Note this type of lower bound readily implies the previously stated worst-case lower bounds known in DP. On the other hand, while the existence of hard distributions follow by the existence of hard datasets (by Yao’s minimax principle), we have chosen to provide explicit constructions of these distributions, just for the sake of clarity. For the pure case, in Lemma A.3 we show that an adaptation of known packing-type lower bounds [23] provides distributional lower bounds for mean estimation, and for the approximate case, we note that fingerprinting distributions induce distributional lower bounds, e.g. [13, 15, 27].
@@
Let $n_{0},t\in\mathbb{N}$ and $\alpha_{0}>0$. Let ${\mathcal{P}}\subseteq{\cal S}_{s}^{t}$, for which there exists a distribution $\mu$ over ${\cal P}^{n_{0}}$ that induces an $(\alpha_{0},\rho_{0})$-distributional lower bound for $(\varepsilon,\delta)$-DP mean estimation. Then, for any $d\geq t$, $n\geq n_{0}$ and $K\leq\min\big{\{}\frac{n}{n_{0}},\frac{d}{t}\big{\}}$, there exists a distribution $\tilde{\mu}$ over $({\cal S}_{s}^{d})^{n}$ that induces a $(\alpha,\rho)$-distributional lower bound for $(\varepsilon,\delta)$-DP mean estimation, where $\alpha\gtrsim\frac{\alpha_{0}n_{0}}{n}\sqrt{\rho_{0}K}$ and $\rho\geq 1-\exp(-\rho_{0}/8)$.
@@
$\mathbb{P}\Big{[}\|{\cal A}(D)-\bar{z}(D)\|_{2}^{2}\geq\big{(}\frac{\alpha_{0}n_{0}}{n}\big{)}^{2}\frac{\rho K}{2}\Big{]}$ $\geq\mathbb{P}\Big{[}\sum_{k=1}^{K}Y_{k}\geq\frac{\rho K}{2}\Big{]}.$
@@
$\mathbb{P}\Big{[}\sum_{k=1}^{K}Y_{k}\geq\frac{\rho K}{2}\Big{]}=\mathbb{P}\Big{[}\sum_{k=1}^{K}\tilde{Y}_{k}\geq\frac{\rho K}{2}\Big{]}\geq\mathbb{P}\Big{[}\sum_{k=1}^{K}W_{k}\geq\frac{\rho K}{2}\Big{]}\geq 1-\exp(-\rho_{0}/8),$
@@
Therefore, $\|{\cal A}(D)-\bar{z}(D)\|_{2}^{2}\gtrsim\Big{(}\frac{\alpha_{0}n_{0}}{n}\Big{)}^{2}\rho K$, with probability $1-\exp(-\rho_{0}/8)$. We conclude that $\tilde{\mu}$ induces a $(\alpha,\rho)$-distributional lower bound for $\varepsilon$-DP mean estimation, as claimed. ∎
The above result needs as input a base lower bound. We note that in the sparse setting, there are packing-based constructions that provide such a base lower bound.
@@
Let $\varepsilon>0$ and $s<d/2$. Then there exists $\mu$ supported on $({\cal S}_{s}^{d})^{n}$ that induces an $(\alpha,\rho)$-distributional lower bound for $\varepsilon$-DP mean estimation with $\alpha\gtrsim\min\left\{1,\frac{s\log(d/s)}{\varepsilon n}\right\}$ and $\rho=1/2$.
@@ 4.2 Lower Bounds for Approximate-DP
Next, we consider the case $\sqrt{s\ln(1/\delta)}/\varepsilon\lesssim n\lesssim\frac{d\sqrt{\ln(1/\delta)}}{\sqrt{s}\varepsilon}$. Let $n_{0}=\sqrt{s\ln(1/\delta)}/\varepsilon$, $t=s$, and $K=\frac{n}{n_{0}}\lesssim\frac{d}{s}$, where the last inequality holds by our regime assumption. The classical $s$-dimensional mean estimation lower bound by [13] provides a $\alpha_{0}\gtrsim 1$ lower bound with constant probability. Hence by Lemma 4.3, the sparse mean estimation problem satisfies a lower bound $\Omega\big{(}\frac{\alpha_{0}n_{0}}{n}\sqrt{K}\big{)}\gtrsim\frac{1}{\sqrt{K}}\gtrsim\frac{[s\ln(1/\delta)]^{1/4}}{\sqrt{\varepsilon n}}$, with constant probability.
@@
5 DP Convex Optimization with Sparse Gradients
We now start with our study of DP convex optimization with sparse gradients. In this section and later, we will impose subsets of the following assumptions:

### v2

@@ 4 Lower Bounds for DP Mean Estimation with Sparse Data
We provide matching lower bounds to those from Section 3. Moreover, although the stated lower bounds are for mean estimation, known reductions imply analogous lower bounds for DP-ERM and DP-SCO [6, 7]. First, for pure-DP we provide a packing-type construction based on sparse vectors. This construction is used in a novel block-diagonal construction, which provides the right low/high-dimensional transition. On the other hand, for approximate-DP, a block diagonal reduction with existing fingerprinting codes [12, 39], suffices to obtain lower bounds that exhibit a nearly tight low/high-dimensional transition. For simplicity, we consider the case of $L=1$, i.e. ${\cal S}_{s}^{d}=\{z\in\mathbb{R}^{d}:\,\|z\|_{0}\leq s,\|z\|_{2}\leq 1\}$; it is easy to see that any lower bound scales linearly in $L$.
@@ 4.1 Lower Bounds for Pure-DP
We say that a probability $\mu$ over ${\cal Z}^{n}$ induces an $(\alpha,\beta)$-distributional lower bound for $(\varepsilon,\delta)$-DP mean estimation if $\inf_{{\cal A}:\,(\varepsilon,\delta)\text{-DP}}\mathbb{P}_{S\sim\mu,{\cal A}}\big{[}\|{\cal A}(S)-\bar{z}(S)\|_{2}\geq\alpha\big{]}\geq\beta.$
Note this type of lower bound readily implies a worst-case lower bound. On the other hand, while the existence of hard distributions follows by the existence of hard datasets (by Yao’s minimax principle), we provide explicit constructions of these distributions, for the sake of clarity. For the pure case, in Lemma A.3 we show that an adaptation of known packing-type lower bounds [23] provides distributional lower bounds for mean estimation, and for the approximate case, we note that fingerprinting distributions induce distributional lower bounds, e.g. [13, 15, 27].
@@
Let $n_{0},t\in\mathbb{N}$. Let $\mu$ be a distribution over $({\cal S}_{s}^{t})^{n_{0}}$ that induces an $(\alpha_{0},\rho_{0})$-distributional lower bound for $(\varepsilon,\delta)$-DP mean estimation. Then, for any $d\geq t$, $n\geq n_{0}$ and $K\leq\min\big{\{}\frac{n}{n_{0}},\frac{d}{t}\big{\}}$, there exists $\tilde{\mu}$ over $({\cal S}_{s}^{d})^{n}$ that induces an $(\alpha,\rho)$-distributional lower bound for $(\varepsilon,\delta)$-DP mean estimation, where $\alpha\gtrsim\frac{\alpha_{0}n_{0}}{n}\sqrt{\rho_{0}K}$ and $\rho\geq 1-\exp(-\rho_{0}/8)$.
@@
$\mathbb{P}\Big{[}\|{\cal A}(D)-\bar{z}(D)\|_{2}^{2}\geq\big{(}\frac{\alpha_{0}n_{0}}{n}\big{)}^{2}\frac{\rho_{0}K}{2}\Big{]}$ $\geq\mathbb{P}\Big{[}\sum_{k=1}^{K}Y_{k}\geq\frac{\rho_{0}K}{2}\Big{]}.$
@@
$\mathbb{P}\Big{[}\sum_{k=1}^{K}Y_{k}\geq\frac{\rho_{0}K}{2}\Big{]}=\mathbb{P}\Big{[}\sum_{k=1}^{K}\tilde{Y}_{k}\geq\frac{\rho_{0}K}{2}\Big{]}\geq\mathbb{P}\Big{[}\sum_{k=1}^{K}W_{k}\geq\frac{\rho_{0}K}{2}\Big{]}\geq 1-\exp(-\rho_{0}/8),$
@@
Therefore, $\|{\cal A}(D)-\bar{z}(D)\|_{2}^{2}\gtrsim\Big{(}\frac{\alpha_{0}n_{0}}{n}\Big{)}^{2}\rho_{0}K$, with probability $1-\exp(-\rho_{0}/8)$. We conclude that $\tilde{\mu}$ induces a $(\alpha,\rho)$-distributional lower bound for $(\varepsilon,\delta)$-DP mean estimation, as claimed. ∎
The above result needs as input a base lower bound. Here, packing-based constructions suffice.
@@
Let $\varepsilon>0$ and $s<d/2$. Then there exists a $(\alpha,\rho)$-distributional lower bound for $\varepsilon$-DP mean estimation over $({\cal S}_{s}^{d})^{n}$ with $\alpha\gtrsim\min\left\{1,\frac{s\log(d/s)}{\varepsilon n}\right\}$ and $\rho=1/2$.
@@ 4.2 Lower Bounds for Approximate-DP
Next, we consider the case $\sqrt{s\ln(1/\delta)}/\varepsilon\lesssim n\lesssim\frac{d\sqrt{\ln(1/\delta)}}{\sqrt{s}\varepsilon}$. Let $n_{0}=\sqrt{s\ln(1/\delta)}/\varepsilon$, $t=s$, and $K=\frac{n}{n_{0}}\lesssim\frac{d}{s}$, where the last inequality holds by our regime assumption. The classic $s$-dimensional mean estimation lower bound by [13] provides a $\alpha_{0}\gtrsim 1$ lower bound with constant probability. Hence by Lemma 4.3, the sparse mean estimation problem satisfies a lower bound $\Omega\big{(}\frac{\alpha_{0}n_{0}}{n}\sqrt{K}\big{)}\gtrsim\frac{1}{\sqrt{K}}\gtrsim\frac{[s\ln(1/\delta)]^{1/4}}{\sqrt{\varepsilon n}}$, with constant probability.
@@
5 Bias-Reduction Method for DP-ERM with Sparse Gradients
We now start with our study of DP-ERM with sparse gradients. In this section and later, we will impose subsets of the following assumptions: